OpenStudy (anonymous):

A taxicab company charges a flat fee of $1.85 plus an additional .40 per quarter mile Write a formula to find the total cost? Use this formula to find the cost for 1 person to travel 8 mi

OpenStudy (ataraxia):

The "flat fee" is a constant, and 0.40 is the ratio wherein it's going to be multiplied by depending on the amount of miles driven. Use the formula: y = mx + b; where "m" is the ratio and "b" the constant. y = 0.40x + 1.85

OpenStudy (ataraxia):

Also, note that y = f(x); and "x" can be anything. For 8 miles, substitute "8" into f(x). f(8) = (8)*0.40 + 1.85 f(8) = 3.2 + 1.85 f(8) = 5.05 For 8 miles, the total cost will be $5.05

OpenStudy (anonymous):

-37?

OpenStudy (anonymous):

Thanks

OpenStudy (ataraxia):

How did you get -37? Haha

OpenStudy (anonymous):

I did a different question